# FOOL: Addressing the Downlink Bottleneck in Satellite Computing with Neural Feature Compression

### v2

@@ 5.4 Compression Model Architecture
Figure 8 illustrates the compression model’s complete architecture. The dashed lines to $Q$ indicate that we either quantize (and subsequently compress) the base latent or the cross-attention weighted latent. For the case of passing $\boldsymbol{\mathrm{y}}_{ca}$ to $Q$, we include an additional residual unit after attention-weighting. We skip applying the attention block to the highest input dimensions to reduce memory and computational costs. The non-linearity between the layers is ReLU to reduce vendor dependency of results from the system performance evaluation in Section 6.4. Residual blocks are two stacked $3\times 3$ convolutions to increase the receptive field with fewer parameters and a residual connection for better gradient flow.
@@
For the remainder of the work, we refer to the compression model as an encoder-decoder pair $enc,dec$. The encoder comprises $g_{a},h_{a},k,f_{ds},c_{a}$, and the entropy coder. The decoder consists of the $g_{s}$ and the entropy decoder. The entropy model $p_{\boldsymbol{\hat{\mathrm{z}}}}(\boldsymbol{\hat{\mathrm{z}}})$ and $h_{s}$ are shared. Note that, despite deploying more components on the constrained sender, the encoder has significantly fewer parameters than the decoder, since we increase the width of the receiver-exclusive components.
@@ 5.6 Image Reconstruction
After training the compression model and freezing encoder weights, FOOL separately trains a reconstruction model that maps $\boldsymbol{\hat{\mathrm{y}}}$ to an approximation $\boldsymbol{\tilde{\mathrm{x}}}$ of the original sample $\boldsymbol{{\mathrm{x}}}$.
@@ 5.6.1 Separate Training over Joint Optimization
We could reduce the distortion $d(\boldsymbol{\mathrm{x}},\boldsymbol{\hat{\mathrm{x}}})$ with a joint objective for training the reconstruction and compression model. While the resulting models would score higher on sum of error benchmarks, the added distortion term will result in higher bitrates. Instead, after optimizing the encoder with the objective of SVBI, we freeze the weights (i.e., “locking” in the rate performance). Then, we leverage the high mutual information between shallow features and the input to recover presentable approximations (Section 3.3.2). Image recovery is closely related to image restoration, such as super-resolution or denoising. The component is exchangeable with the state-of-the-art, as it is orthogonal to the compression task. For this work, we select SwinIR [55] due to its relative recency, computational efficiency, and simplicity.
@@ 5.6.2 Reconstruction Does not Replace Decoders
Approximations from decoders (Section 5.5) resulting in (near) lossless prediction, would evidence $\hat{\boldsymbol{\mathrm{y}}}$ has sufficient information to reconstruct a sample for the input layers that result in comparable task performance. Hence, after training the encoder, we could replace all decoders with the reconstruction model to approximate the input sample. Nevertheless, sufficiency may not directly result in lossless prediction since artifacts perturb the distribution of samples derived from image recognition. We could account for the perturbation by finetuning for a relatively small number of iterations (indirectly shown in [15, 17]). The downside is that operators must maintain, store, and serve different versions of the otherwise identical backbones for each client separately. Worse, they may need to re-train the predictors of the various downstream tasks for each backbone. Instead, we can train small decoders that directly map the low-dimensional encoder output to an adequate representation (i.e., the input to deeper layers). There are two advantages to introducing multiple decoders over multiple backbone weights. First, the number of additional weights operators must maintain only scales with supported backbones and not the number of backbone-task pairs. Second, the small decoder weights incur considerably less training and storage overhead than the weights of massive backbones.
@@ 5.7.1 Rate-Distortion Loss Function for SVBI
To simplify the loss expression, we treat the components related to keypoints as part of $h_{a}$ if we exclusively use it as a hint for the side-information network. Alternatively, as part of $g_{a}$, when used as an additional operation to reduce spatial dependencies before quantizing and entropy coding the latent. Analogous to the SVBI training objective [17, 15], we have a parametric analysis transform $g_{a}(x;\theta)$ that maps $x$ to a latent vector $z$. Then, a quantizer $Q$ discretizes $z$ to $\hat{z}$ for lossless entropy coding. Since we rely on HD (Figure 2) as a distortion function, the parametric synthesis transforms $g_{s}(\hat{x};\phi)$ that maps $\hat{y}$ to an approximation of a representation $\tilde{h}$. As introduced in [42], we apply uniform quantization $Q$, but replace $Q$ with continuous relaxation by adding uniform noise $\eta\thicksim\mathcal{U}(-\frac{1}{2},\frac{1}{2})$ during training for gradient computation.
Without a hyperprior, the loss function is:
@@ 5.7.2 Mapping Encoder Output to Target Representations
After training the first $enc,dec_{1}$ pair, FOOL freezes $enc$ weights, i.e., only applying the distortion term of the loss in Equation 19, for subsequent decoders $dec_{2},dec_{3}\dots dec_{n}$ (Section 5.5). Lastly, FOOL treats the reconstruction model $rec$ as a decoder with the identity function as a teacher, i.e., $\mathcal{H}_{rec}(x)=x$. The difference from other decoders is that the target representation must be human-interpretable. Hence, we draw from image restoration that primarily focuses on synthesizing high-quality images and replace $\operatorname{sse}$ with the Charbonnier loss [56]:
@@
It is out of this work’s scope to exhaustively evaluate image restoration methods. Rather, the focus is to provide empirical evidence for the claims in Section 3.3.1. We simply found that despite performing comparable to other sums of error losses on benchmark metrics, subjective visual inspections revealed that Charbonnier resulted in smoother images.
@@ 6.1.1 Testbed
6.1.2 Third-Party Detection Models
@@
While the work in [17] did not explicitly evaluate object detection tasks, the support for two-stage detectors follows from the codec sufficiently approximating the representation of the feature extractor (i.e., the first stage). However, it is not apparent whether the general SVBI framework yields gains over image codecs when the targets are one-stage detectors. Therefore, to replicate a representative service for inference on aerial or satellite imagery, we apply simple transfer learning on open-source weights [61] for YOLOv5 and YOLOv8.
Image codecs pass a sample $dec(enc(\boldsymbol{\mathrm{x}}))=\boldsymbol{\hat{\mathrm{x}}}$ to the input layers of a target model. Feature codecs (i.e., SVBI methods) skip the shallow layers and pass $dec(enc(\boldsymbol{\mathrm{x}}))=\boldsymbol{\hat{\mathrm{h}}}$ to the deeper layers. Detection models with the same architecture share the frozen shallow layers (i.e., layers until the first non-residual connection).
We select detection tasks to represent varying mission conditions. DOTA-2 [62] for a more coarse-grained aerial task with comparatively lower Ground Sample Distance (GSD) and larger objects. SpaceNet-3 [63] for urban tasks (e.g., for traffic control) with high image resolutions. Lastly, xView [64] for disaster response systems where detection models rely on fine-grained details. To simplify the already intricate evaluation setup and to ease reproducibility, we deliberately refrain from more refined transfer learning methods. We merely require detection models with mAP scores that are moderately high to determine whether a codec can preserve fine-grained details for EO tasks on satellite imagery. For each architecture, we jointly finetune the deeper layers and train separate predictors that achieve around 35-65% mAP@50.
@@ 6.3 Image Reconstruction Quality
Figure 13 further visualizes the potential of reliably recovering fine-grained from the compressed latent space. Naturally, it should be possible to finetune the TinyLic-MQ to improve perceptual quality analogous to FOOl-FT. However, TinyLIC is still a significantly costlier model, with a worse rate and prediction performance. More pressingly, we stress that the reliability of a restoration model is bound by the available signals in the compressed latent space. Accordingly, we deliberately avoid generative models that prioritize realism over structural integrity. In particular, prioritizing realism over reliability defeats the primary purpose of image restoration, i.e., intervention by human experts in critical EO applications. A model outperforming experts does not imply that predictions may inexplicably be false. Specifically, where human cost is involved (e.g., disaster warning or relief [5]) it is paramount that experts can trust the codec to not include extrapolated elements to an image.
@@ 6.4.1 Processing Throughput and Transfer Cost Reduction
Table VI summarizes the configuration that maximizes profiler selection by TCR/s for all models on each device separately. Since bitrate variance is low between DOTA-2, SpaceNet-3, and xView, we average the bpp (Section 6.2) on the validation sets. The bold Model value indicates the adequate size of each model on a device, i.e., the model we will deploy to measure data volume downlinking in the following experiments. The bold TCR/s marks the highest overall value for a device, i.e., we can expect applying FOOL over BSVBI to result in considerably more downlinkable data on all devices. However, due to keypoint extraction and the ITA layers, FOOL’s processing throughput is slower than BVSBI’s. The overhead is particularly punishing for the most constrained device (i.e., the previous-generation TX2), where BSVBI-M has slightly higher TCR/s than FOOL-M, despite the latter’s significantly better compression performance. Moreover, the profiler selects FOOL-S over FOOL-M/-L for the low-end current generation Orin Nano and high-end last-gen past generation NX. Conversely, the profiler decides on the mid-sized model for BSVBI across all devices despite FOOL’s compression performance scaling better.
@@ 6.4.2 Model Inference with Concurrent Task Execution
Table VII summarizes the results from running the entire compression pipeline with concurrent task execution using the configurations that maximize TCR/s from Table VI. The bold values in the TCR/s dec. column indicates the size with the highest decrease. A file includes all model artifacts output by the neural codec’s DNN components for a single tile, i.e., the pipeline still needs to entropy code them to match the bpp in TCR/s calculations. File size refers to the storage requirements per tile of the encoder output tensors, i.e., the data volume the rANS process encodes. We compute file size by a worst-case upper bound by the encoder output tensor dimensionality (Section 5.4) without serialization formats that could exploit the sparsity of $\boldsymbol{\hat{\mathrm{y}}}$ and $\boldsymbol{\hat{\mathrm{z}}}$. There are two essential findings from the results. First, the rANS process can consume tasks considerably faster than the inference process can produce them, i.e., there is no risk of backpressure within the pipeline. Second, there is only a minimal percentage decrease in TCR/s across all devices and models relative to sequential execution. Hence, we argue that the pipeline successfully offsets the runtime overhead as claimed in Section 4.3 even without relying on a precomputed lookup table (e.g., tANS in ZSTD [77]).
The results are unsurprising when viewing the CPU (red) and GPU (blue) load of DNN inference without CPU-bound concurrent tasks in Figure 15.
@@ 6.4.3 Downlinkable Data Volume
Notice the logarithmic scale, i.e., FOOL improves downlinking using bent pipes by over two orders of magnitude without relying on prior information on the downstream tasks or crude filtering methods. For example, given Maxar’s WorldView-3 conditions [29], it would be possible to downlink roughly 9TB of sensor data per pass, before reaching downlink saturation. As a comparison, the state-of-the-art filtering method in [27] reports a 3$\times$ improvement, based on a definition of value. Note that to provide a realistic presentation of the opportunities SVBI provides, we assume that a nanosatellite processes tiles until reaching a downlink segment. Moreover, we disregard the “computational deadline”, i.e., it can process all the data before reaching a ground segment. This is reasonable since there should always be enough data to process. If not produced by a single onboard sensor, constellations may designate certain satellites as compression nodes using reliable, high-capacity local communication channels [78]. Further, from the throughput evaluation and the orbital period, it is inferrable that even the low-end current-generation Orin Nano would barely miss the computational deadline. We remind the reader that the presented throughput measures represent a worst-case lower bound, as no vendor-specific software or hardware optimization was applied.
@@ 7 Discussions and Limitations
The following elaborates on limitations and highlights some research directions we find particularly promising to increase mission value. A general recommendation is to leverage recent advancements in image restoration [79] by utilizing generative models as robust priors and ensuring enough signals are available in the downlinked data to preserve structures at salient regions.
@@ 7.1 Downlink Saturation Handling
Readers may notice the omission of handling downlink saturation A simple solution is to apply the compression model after the initial filtering. However, this contradicts FOOL’s general design philosophy that it should not limit practicability due to prior knowledge of downstream tasks, excluding filtering as a preprocessing step, such as data deduplication. Therefore, we propose future work should prioritize segments according to their entropy in latent space and apply inpainting methods [80]. Intuitively, the lower the remaining uncertainty, given other tiles, the more reliable models can predict missing tiles.
@@ 7.2 Reliance on Foundational Models
Earth Observation (EO) requires considerations not included in common object detection objectives and architectural components, and widespread foundational models for satellite imagery have yet to emerge. Nonetheless, proprietary offerings already exist [81], and we argue that the community drives to open-source solutions will inevitably mitigate the limitation. We worked around not having access to an EO-native foundational model by only freezing the shallow layers to train the predictors. In other words, each predictor is complemented with a large backbone trained for a certain GSD range, i.e., not preserving the details with adequate granularity would result in predictive loss.
@@ Acknowledgment
We thank Alexander Knoll for providing us with the hardware infrastructure. Kerstin Bunte for her valuable suggestions. Florian Kowarsch for the fruitful discussions we held.

### v3

@@ 5.4 Compression Model Architecture
Figure 8 illustrates the compression model’s complete architecture. The dashed lines to $Q$ indicate that we either quantize (and subsequently compress) the base latent or the cross-attention weighted latent.
@@
For the case of passing $\boldsymbol{\mathrm{y}}_{ca}$ to $Q$, we include an additional residual unit after attention-weighting. We skip applying the attention block to the highest input dimensions to reduce memory and computational costs. The non-linearity between the layers is ReLU to reduce vendor dependency of results from the system performance evaluation in Section 6.4. Residual blocks are two stacked $3\times 3$ convolutions to increase the receptive field with fewer parameters and a residual connection for better gradient flow. For the remainder of the work, we refer to the compression model as an encoder-decoder pair $enc,dec$. The encoder comprises $g_{a},h_{a},k,f_{ds},c_{a}$, and the entropy coder. The decoder consists of the $g_{s}$ and the entropy decoder. The entropy model $p_{\boldsymbol{\hat{\mathrm{z}}}}(\boldsymbol{\hat{\mathrm{z}}})$ and $h_{s}$ are shared. Note that, despite deploying more components on the constrained sender, the encoder has significantly fewer parameters than the decoder since we increase the width of the receiver-exclusive components.
@@ 5.6 Image Reconstruction
FOOL trains the compression and image reconstruction models in two stages. After training the compression model and freezing encoder weights, it separately trains a reconstruction model that maps $\boldsymbol{\hat{\mathrm{y}}}$ to an approximation $\boldsymbol{\tilde{\mathrm{x}}}$ of the original sample $\boldsymbol{{\mathrm{x}}}$.
@@ 5.6.1 Separate Training over Joint Optimization
We could reduce the distortion $d(\boldsymbol{\mathrm{x}},\boldsymbol{\hat{\mathrm{x}}})$ with a joint objective for training the reconstruction and compression model. While the resulting models would score higher on the sum of error benchmarks, the added distortion term will result in higher bitrates. Instead, after optimizing the encoder with the objective of SVBI, we freeze the weights (i.e., “locking” in the rate performance). Then, we leverage the high mutual information between shallow features and the input to recover presentable approximations (Section 3.3.2).
Image recovery is closely related to image restoration, such as super-resolution or denoising. The component is exchangeable with the state-of-the-art, as it is orthogonal to the compression task. For this work, we select SwinIR [55] due to its relative recency, computational efficiency, and simplicity.
@@ 5.6.2 Reconstruction Does not Replace Decoders
Approximations from decoders (Section 5.5) resulting in (near) lossless prediction would evidence $\hat{\boldsymbol{\mathrm{y}}}$ has sufficient information to reconstruct a sample for the input layers that result in comparable task performance. Hence, after training the encoder, we could replace all decoders with the reconstruction model to approximate the input sample. Nevertheless, sufficiency may not directly result in lossless prediction since artifacts perturb the reconstructed samples.
We could account for the perturbation by finetuning for a relatively small number of iterations [17]. The downside is that operators must maintain, store, and serve different versions of the otherwise identical backbones for each client separately. Worse, they may need to re-train the predictors of the various downstream tasks for each backbone. Instead, we train small decoders that directly map the low-dimensional encoder output to an adequate representation, i.e., FOOL does not pass the reconstruction to prediction models for downstream tasks. There are two advantages to introducing multiple decoders over multiple backbone weights. First, the number of additional weights operators must maintain only scales with supported backbones and not the number of backbone-task pairs. Second, the small decoder weights incur considerably less training and storage overhead than the weights of massive backbones.
@@ 5.7.1 Rate-Distortion Loss Function for SVBI
To simplify the loss expression, we treat the components related to keypoints as part of $h_{a}$ if we exclusively use it as a hint for the side-information network. Alternatively, it may be used as the final block of $g_{a}$ before quantizing and entropy coding the latent. Analogous to the SVBI training objective [17, 15], we have a parametric analysis transform $g_{a}(x;\theta)$ that maps $x$ to a latent vector $z$. Then, a quantizer $Q$ discretizes $z$ to $\hat{z}$ for lossless entropy coding. Since we rely on HD (Figure 2) as a distortion function, the parametric synthesis transforms $g_{s}(\hat{x};\phi)$ that maps $\hat{y}$ to an approximation of a representation $\tilde{h}$. As introduced in [42], we apply uniform quantization $Q$, but replace $Q$ with continuous relaxation by adding uniform noise $\eta\thicksim\mathcal{U}(-\frac{1}{2},\frac{1}{2})$ during training for gradient computation.
Without a hyperprior, the loss is:
@@ 5.7.2 Mapping Encoder Output to Target Representations
After training the first $enc,dec_{1}$ pair, FOOL freezes $enc$ weights, i.e., only applying the distortion term of the loss in Equation 19, for subsequent decoders $dec_{2},dec_{3}\dots dec_{n}$ (Section 5.5). Lastly, FOOL treats the reconstruction model $rec$ as a decoder and assigns the identity function as its teacher, i.e., $\mathcal{H}_{rec}(x)=x$. Unlike for other decoders, the target representation must be human-interpretable. Hence, we train the decoder for image reconstruction using the Charbonnier Loss [56]
@@
where $\epsilon$ is a small constant we set as $2\cdot 10^{-3}$. It is out of this work’s scope to exhaustively evaluate image restoration methods. Rather, the focus is to provide empirical evidence for the claims in Section 3.3.1. We simply found that despite performing comparable to other sums of error losses on benchmark metrics, using Charbonnier results in more stable training.
@@ 6.1.1 Testbed
6.1.2 Third-Party Detection Models & Target Tasks
@@
While the work in [17] did not explicitly evaluate object detection tasks, the support for two-stage detectors follows from the codec sufficiently approximating the representation of the feature extractor (i.e., the first stage). However, it is not apparent whether the general SVBI framework yields gains over image codecs when the targets are one-stage detectors. Therefore, to replicate a representative service for inference on aerial or satellite imagery, we apply simple transfer learning on open-source weights [61] for YOLOv5 and YOLOv8. Image codecs pass a sample $dec(enc(\boldsymbol{\mathrm{x}}))=\boldsymbol{\hat{\mathrm{x}}}$ to the input layers of a target model. Feature codecs (i.e., SVBI methods) skip the shallow layers and pass $dec(enc(\boldsymbol{\mathrm{x}}))=\boldsymbol{\hat{\mathrm{h}}}$ to the deeper layers. Detection models with the same architecture share the frozen shallow layers (i.e., layers until the first non-residual connection). We associate one task as outlined by the test labels for each of the three dataset separately. The tasks represent varying mission conditions. DOTA-2 [62] for a more coarse-grained aerial task with comparatively lower Ground Sample Distance (GSD) and larger objects. SpaceNet-3 [63] for urban tasks (e.g., for traffic control) with high image resolutions. Lastly, xView [64] for disaster response systems where detection models rely on fine-grained details. Lastly, image reconstruction is treated distinctly as single task for the reconstruction model, and not the detection pipeline, by combining the images from the three test sets.
To simplify the already intricate evaluation setup and to ease reproducibility, we deliberately refrain from more refined transfer learning methods. We merely require detection models with mAP scores that are moderately high to determine whether a codec can preserve fine-grained details for EO tasks on satellite imagery. For each architecture, we jointly finetune the deeper layers and train separate predictors that achieve around 35-65% mAP@50.
@@ 6.3 Image Reconstruction Quality
Figure 13 further visualizes the potential of reliably recovering fine-grained from the compressed latent space. Naturally, it should be possible to finetune the TinyLic-MQ to improve perceptual quality analogous to FOOl-FT. However, TinyLIC is still a significantly costlier model, with a worse rate and prediction performance. More pressingly, we stress that the reliability of a restoration model is bound by the available signals in the compressed latent space. Accordingly, we deliberately avoid generative models that prioritize realism over structural integrity. Prioritizing realism over reliability defeats the primary purpose of image restoration, i.e., intervention by human experts in critical EO applications. A model outperforming experts does not imply that predictions may inexplicably be false. In particular, where human cost is involved (e.g., disaster warning or relief [5]) it is paramount that experts can trust the codec to not include extrapolated elements to an image.
@@ 6.4.1 Processing Throughput and Transfer Cost Reduction
Table VI summarizes the configuration that maximizes profiler selection by TCR/s for all models on each device separately. Since bitrate variance is low between DOTA-2, SpaceNet-3, and xView, we average the bpp (Section 6.2) on the validation sets. The bold Model value indicates the adequate size of each model family on a device, i.e., the model we will deploy to measure data volume downlinking in the following experiments. The bold TCR/s marks the highest overall value for a device, i.e., we can expect applying FOOL over BSVBI to result in considerably more downlinkable data on all devices. However, due to keypoint extraction and the ITA layers, FOOL’s processing throughput is slower than that of BVSBI. The overhead is particularly punishing for the most constrained device (i.e., the previous-generation TX2), where BSVBI-M has slightly higher TCR/s than FOOL-M despite the latter’s significantly better compression performance. Moreover, the profiler selects FOOL-S over FOOL-M/-L for the low-end current generation Orin Nano and high-end last-gen past generation NX. Conversely, the profiler decides on the mid-sized model for BSVBI across all devices despite FOOL’s compression performance scaling better.
@@ 6.4.2 Model Inference with Concurrent Task Execution
Table VII summarizes the results from running the entire compression pipeline with concurrent task execution using the configurations that maximize TCR/s from Table VI. The bold values in the TCR/s dec. column indicates the size with the highest decrease. A file includes all model artifacts output by the neural codec’s DNN components for a single tile, i.e., the pipeline still needs to entropy code them to match the bpp in TCR/s calculations. File size refers to the storage requirements per tile of the encoder output tensors, i.e., the data volume the rANS process encodes. We compute file size by a worst-case upper bound by the encoder output tensor dimensionality (Section 5.4) without serialization formats that could exploit the sparsity of $\boldsymbol{\hat{\mathrm{y}}}$ and $\boldsymbol{\hat{\mathrm{z}}}$. There are two essential findings from the results. First, the rANS process can consume tasks considerably faster than the inference process can produce them, i.e., there is no risk of backpressure within the pipeline. Second, there is only a minimal percentage decrease in TCR/s across all devices and models relative to sequential execution. Hence, we argue that the pipeline successfully offsets the runtime overhead as claimed in Section 4.3 even without relying on a precomputed lookup table (e.g., tANS in ZSTD [77]). The results are unsurprising when viewing the CPU and GPU load of DNN inference without CPU-bound concurrent tasks in Figure 15.
@@ 6.4.3 Downlinkable Data Volume
Notice the logarithmic scale, i.e., FOOL improves downlinking using bent pipes by over two orders of magnitude without relying on prior information on the downstream tasks or crude filtering methods. For example, given Maxar’s WorldView-3 conditions [29], it would be possible to downlink roughly 9TB of sensor data per pass before reaching downlink saturation. As a comparison, the state-of-the-art filtering method in [27] reports a 3$\times$ improvement based on a definition of value. Note that to provide a realistic presentation of the opportunities SVBI provides, we assume that a nanosatellite processes tiles until reaching a downlink segment. Moreover, we disregard the “computational deadline”, i.e., it can process all the data before reaching a ground segment. This is reasonable since there should always be enough data to process. If not produced by a single sensor, constellations may designate certain satellites as compression nodes using reliable, high-capacity local communication channels [78]. Further, it is inferrable that even the low-end current-generation Orin Nano without any vendor-specific optimization would barely miss the computational deadline.
@@ 7.1 Downlink Saturation Handling
This work has omitted to handle downlink saturation, i.e., prioritizing salient data when compression cannot sufficiently reduce the volume in time. Existing work suggests to apply intelligent adhoc filtering. However, as argued in Section 2, such filtering relies on strong assumptions that bias the downlinked data towards a small subset of tasks clients may be interested in. Therefore, we suggest that filtering should prioritize tiles that provide sufficient information to recover filtered tiles with a generative model. Intuitively, the lower the remaining uncertainty, given other tiles, the more reliable generative models can recover missing tiles.
@@ 7.2 Reliance on Foundational Models
Earth Observation (EO) requires considerations not included in common object detection objectives and architectural components, and widespread foundational models for satellite imagery have yet to emerge. Nonetheless, proprietary offerings already exist [79], and we argue that the community drives to open-source solutions will inevitably mitigate the limitation. We worked around not having access to an EO-native foundational model by only freezing the shallow layers to train the predictors, i.e, each predictor is complemented with a backbone suitable for a particular sensor configuration.
@@ Acknowledgment
We thank Alexander Knoll for providing us with the hardware infrastructure. Kerstin Bunte for her valuable suggestions. Florian Kowarsch for the fruitful discussions we held. The authors acknowledge TU Wien Bibliothek for financial support through its Open Access Funding Programme.